# Relational Conformal Prediction for Correlated Time Series

### v1

@@ Appendix E Evaluation metrics
Appendix F Hyperparameters and empirical setup
@@
For CoRel we tuned the number of neurons in the STGNN with a small grid search on $10\%$ of the calibration data. For the experiments on real-word data, the model was training for a maximum of $100$ epochs the calibration set. Each epoch consisted of a maximum of $50$ mini-batches of size $64$. We used the Adam optimizer [80] with an initial learning rate of $0.003$ and reduced by $75\%$ every $20$ epochs. We used a fixed number of $K=20$ neighbors for the grapn learning module. We used the same hyperparameters and model selection procedure for CoRNN.
@@
For HopCPT we followed the model selection procudere described in [21]. The model was trained for $3000$ epochs using a batch size of $4$ time series. We adopted the paper’s AdamW optimizer configuration with standard parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=0.01$) and tuned the model by running the same hyperparameter configurations searched by [21]. All the remaining hyperparameters were set accordingly to the original paper.
@@
SCPI
As already mentioned, we used the SCPI [19] implementation provided by Auer et al. [21] and followed an analogous protocol for training the mode. In particular, SCPI was run using a fixed window length of $100$ time steps for all experiments, corresponding to the longest setting in the original paper. The computational demands of SPCI were substantial, as separate quantile random forests had to be trained for each combination of node and target coverage level, making extensive hyperparameter tuning impractical. To further manage the computational costs, we trained each SPCI model only once on the calibration data, rather than implementing the time-adaptive approach where models are re-trained as new observations become available.
Appendix G Optimization of the PI width
Here, we show how the width of the PI can be reduced by searching for an appropriate quantile offset $\beta$. Specifically, we perform the optimization described in Eq. 16, which identifies pairs of quantiles yielding the same coverage $1-\alpha$ while achieving the smallest interval width. We summarize our results in Tab. 4. As we can see, the procedure consistently finds intervals with a smaller width while retaining approximately the same coverage, whether using CoRel or the CoRNN variant.
@@ Appendix G Optimization of the PI width
METR-LA CoRNN -1.52 $\pm$ 0.35 19.26 $\pm$ 0.27 -1.01 $\pm$ 0.28 19.45 $\pm$ 0.13 -0.40 $\pm$ 0.57 16.23 $\pm$ 0.36
$\beta$ optim. -1.50 $\pm$ 0.32 17.17 $\pm$ 0.21 -1.28 $\pm$ 0.18 17.16 $\pm$ 0.17 -0.64 $\pm$ 0.44 14.95 $\pm$ 0.32
CoRel -1.15 $\pm$ 0.48 15.67 $\pm$ 0.19 -0.47 $\pm$ 0.33 16.14 $\pm$ 0.38 -0.80 $\pm$ 0.49 15.73 $\pm$ 0.37
$\beta$ optim. -1.40 $\pm$ 0.52 14.58 $\pm$ 0.15 -0.91 $\pm$ 0.43 15.00 $\pm$ 0.33 -0.89 $\pm$ 0.48 14.82 $\pm$ 0.39
CER-E CoRNN -2.12 $\pm$ 0.19 2.27 $\pm$ 0.01 -2.05 $\pm$ 0.12 2.25 $\pm$ 0.01 -2.19 $\pm$ 0.13 2.04 $\pm$ 0.02
$\beta$ optim. -2.41 $\pm$ 0.15 2.20 $\pm$ 0.01 -2.31 $\pm$ 0.08 2.19 $\pm$ 0.01 -2.47 $\pm$ 0.13 2.00 $\pm$ 0.02
CoRel -3.56 $\pm$ 0.24 1.92 $\pm$ 0.01 -3.10 $\pm$ 0.33 1.95 $\pm$ 0.01 -4.90 $\pm$ 0.34 1.82 $\pm$ 0.03
$\beta$ optim. -4.09 $\pm$ 0.27 1.85 $\pm$ 0.01 -3.38 $\pm$ 0.32 1.87 $\pm$ 0.01 -5.28 $\pm$ 0.36 1.75 $\pm$ 0.03
AQI CoRNN -0.42 $\pm$ 0.84 75.85 $\pm$ 2.13 -0.10 $\pm$ 0.46 77.37 $\pm$ 1.32 -0.17 $\pm$ 0.42 $7$$\pm$ 0.43
$\beta$ optim. -0.36 $\pm$ 0.85 72.96 $\pm$ 1.96 0.05 $\pm$ 0.42 74.73 $\pm$ 1.16 -0.12 $\pm$ 0.40 75.58 $\pm$ 0.43
CoRel -1.53 $\pm$ 1.60 70.47 $\pm$ 3.60 -2.93 $\pm$ 1.76 71.04 $\pm$ 4.76 -2.50 $\pm$ 1.12 71.88 $\pm$ 2.87
$\beta$ optim. -2.78 $\pm$ 1.79 67.30 $\pm$ 3.37 -3.70 $\pm$ 1.79 67.95 $\pm$ 4.31 -3.51 $\pm$ 1.04 69.13 $\pm$ 2.73

### v2

@@ Appendix E Evaluation metrics
Appendix F Hyperparameters and experimental setup
@@
For CoRel we tuned the number of neurons in the STGNN with a small grid search on $10\%$ of the calibration data. We used the same model selection procedure for CoRNN but also tuned the number of GRU layers. For the experiments on real-world data, the model was trained for a maximum of $100$ epochs on the calibration set. Each epoch consisted of a maximum of $50$ mini-batches of size $64$. We used the Adam optimizer [Kingma and Ba, 2015] with an initial learning rate of $0.003$ and reduced by $75\%$ every $20$ epochs. We used a fixed number of $K=20$ neighbors for the graph learning module. For the GPVAR experiment in Sec. 5.2, we fixed the number of neurons to $16$ for each layer and used an embedding size of $8$. For the adaptation experiment in Sec. 5.3, we use a reference configuration with $64$ neurons in each encoder/decoder layers and an embedding size of $32$. We train the entire model on the full calibration set and use adaptation at test time. In particular, at test time, we fine-tune the node embeddings every $M$ time steps, where $M$ corresponds to $\frac{1}{6}$ of the test set length. For fine-tuning, we fit the embeddings by running $25$ epochs of maximum $10$ batches each by using samples from the last $M$ steps with a fixed learning rate of $0.001$.
@@
For HopCPT we followed the model selection procudere described in [Auer et al., 2023]. The model was trained for $3000$ epochs using a batch size of $4$ time series. We adopted the paper’s AdamW optimizer configuration with standard parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=0.01$) and tuned the model by running the same hyperparameter configurations searched by [Auer et al., 2023]. All the remaining hyperparameters were set accordingly to the original paper.
@@
SPCI
As already mentioned, we used the SPCI [Xu and Xie, 2023b] implementation provided by Auer et al. [2023] and followed an analogous protocol for training the mode. In particular, SPCI was run using a fixed window length of $100$ time steps for all experiments, corresponding to the longest setting in the original paper. The computational demands of SPCI were substantial, as separate quantile random forests had to be trained for each combination of node and target coverage level, making extensive hyperparameter tuning impractical. To further manage the computational costs, we trained each SPCI model only once on the calibration data, rather than implementing the time-adaptive approach where models are re-trained as new observations become available.
Appendix G Additional results
Figure 2: Comparison between (a) the true graph and (b) the graph structure learned by CoRel in GPVAR.
G.1 Qualitative analysis of the learned graph on GPVAR
Fig. 2 provides a comparison of the graph learned on GPVAR by CoRel in Sec. 5.2. In particular, the rightmost figure is obtained by taking the top-$K$ scores associated to each row [^2] The figure does not show edge scores associated to dummy nodes; see App. D. The figure shows that the learned graph includes true edges plus additional links. Note that perfectly recovering the ground truth graph is not required to achieve optimal performance here.
G.2 Optimization of the PI width
Here, we show how the width of the PI can be reduced by searching for an appropriate quantile offset $\beta$. Specifically, we perform the optimization described in Eq. 16, which identifies pairs of quantiles yielding the same coverage $1-\alpha$ while achieving the smallest interval width. We summarize our results in Tab. 4. As we can see, the procedure consistently finds intervals with a smaller width at the price of slightly reducing the coverage, whether using CoRel or the CoRNN variant.
@@ G.2 Optimization of the PI width
METR-LA CoRNN -0.70 $\pm$ 0.20 19.69 $\pm$ 0.27 -0.52 $\pm$ 0.56 19.07 $\pm$ 0.42 -0.67 $\pm$ 0.47 16.02 $\pm$ 0.13
$\beta$ optim. -0.80 $\pm$ 0.28 17.71 $\pm$ 0.28 -0.96 $\pm$ 0.35 17.16 $\pm$ 0.46 -0.88 $\pm$ 0.44 14.98 $\pm$ 0.17
CoRel -1.25 $\pm$ 0.80 14.24 $\pm$ 0.53 -0.97 $\pm$ 0.51 14.43 $\pm$ 0.38 -1.31 $\pm$ 0.48 14.36 $\pm$ 0.20
$\beta$ optim. -1.49 $\pm$ 0.56 13.23 $\pm$ 0.46 -1.26 $\pm$ 0.42 13.47 $\pm$ 0.35 -1.44 $\pm$ 0.45 13.46 $\pm$ 0.20
CER-E CoRNN -2.09 $\pm$ 0.56 1.93 $\pm$ 0.01 -1.70 $\pm$ 0.37 1.99 $\pm$ 0.02 -2.45 $\pm$ 0.41 1.85 $\pm$ 0.03
$\beta$ optim. -2.63 $\pm$ 0.65 1.82 $\pm$ 0.01 -2.28 $\pm$ 0.31 1.88 $\pm$ 0.02 -3.10 $\pm$ 0.52 1.78 $\pm$ 0.03
CoRel -3.86 $\pm$ 0.53 1.83 $\pm$ 0.02 -3.81 $\pm$ 0.30 1.83 $\pm$ 0.02 -4.87 $\pm$ 0.51 1.79 $\pm$ 0.04
$\beta$ optim. -4.16 $\pm$ 0.59 1.75 $\pm$ 0.02 -4.10 $\pm$ 0.21 1.75 $\pm$ 0.02 -5.13 $\pm$ 0.51 1.72 $\pm$ 0.03
AQI CoRNN -1.45 $\pm$ 0.61 73.88 $\pm$ 1.17 -1.08 $\pm$ 0.22 74.02 $\pm$ 0.74 -1.50 $\pm$ 0.63 70.42 $\pm$ 0.94
$\beta$ optim. -1.90 $\pm$ 0.54 70.17 $\pm$ 1.22 -1.10 $\pm$ 0.25 70.95 $\pm$ 0.68 -1.38 $\pm$ 0.62 68.24 $\pm$ 0.85
CoRel -2.40 $\pm$ 1.97 70.44 $\pm$ 2.69 -3.54 $\pm$ 1.80 68.31 $\pm$ 4.49 -2.61 $\pm$ 1.92 67.76 $\pm$ 3.12
$\beta$ optim. -2.83 $\pm$ 1.86 67.35 $\pm$ 2.31 -4.19 $\pm$ 1.96 65.20 $\pm$ 4.20 -3.22 $\pm$ 2.45 65.32 $\pm$ 3.43